Two consecutive numbers have a sum of 35 and a product of 306. Find the smaller number.
Found 2 solutions by Theo, nihar*2013:
Answer by Theo(13342)

x is the first number
x+1 is the second number.

you get:

x + x + 1 = 35
x * (x+1) = 306

simplify to get:

2x+1 = 35
x^2+x = 306

from the first equation, solve for x to get x = 17.

it looks like one of your numbers has to be 17.

then the other number must be 18 because the numbers are consecutive starting from 17.

multiply 17 * 18 and you get 306 so this confirms that the numbers are 17 and 18.

if you had solved for x in the second eqaution, it would have given you x = 17 as well.

x^2 + x = 306 is the second equation.
subtract 306 from both sides of the equation to get x^2 + x - 306 = 0
factor the equation to get (x+18) * (x-17) = 306
solve for x to get x = -18 or x = 17

x can't be -18 because -18 * -17 results in -306 and not 306.
therefore x has to be 17.

17*18 = 306
17+18 = 35

since you are asked to find the smaller number, then your solution is that the smaller number is equal to 17.






Answer by nihar*2013(34)

.
try my free science app
http://hiranituitions.com/category/study-material/
Try my sci app how to download and install is mentioned on above link Rate it share and help me to make education world wide that app has Who wants to be billionaire like and it has many topic of science
Physics
1. An introduction to Nano technology
2. Light Reflection and Refraction
3. Dispersion of light and Natural optical phenomena
4. Electricity
5. Magnetic effect of electric current
6. Universe
Chemistry
1. Acid Base and Salt
2. Metals
3. Non Metals
4. Mineral Coal and Mineral oil
5. Organic compounds
Bio logy
1. Nutrition and Respiration
2. Transportation circulation and Excretion in Organisms
3. Control and coordination In organisms
4. Reproductions in organisms
5. Heredity and Evolution
6. Our Environment
Download and enjoy + Enjoy this solution thx and yes Give positive comment on my site also